OpenStudy (anonymous):

What is the value of x? 6(-3-x)-2x=14

pooja195 (pooja195):

(6)(−3)+(6)(−x)+−2x=14(Distribute)

OpenStudy (anonymous):

so how do get rid of the x

pooja195 (pooja195):

(6)(−3)+(6)(−x)+−2x=14(Distribute) −18+−6x+−2x=14 (−6x+−2x)+(−18)=14(Combine Like Terms)

pooja195 (pooja195):

Whats -6-(+2) ?

OpenStudy (anonymous):

-8

pooja195 (pooja195):

−8x+−18=14 from here solve for x

OpenStudy (anonymous):

x=4

pooja195 (pooja195):

x=-4 but yes you have the right idea :)

OpenStudy (anonymous):

ty